OpenStudy (errinss):

A cylinder is full at 471 cm3 and has a radius of 5 cm. It currently contains 314 cm3 of water. What will the difference in the height of the water be when the cylinder is full? Use 3.14 to approximate pi.

OpenStudy (errinss):

@Hauntedwoodsgal do you know how to do this?

OpenStudy (thatonegirl_):

So you are trying to solve for the height for when the cylinder is completely full and when it is at 471cm3 correct?

OpenStudy (thatonegirl_):

314*

OpenStudy (errinss):

ya

OpenStudy (eskelle):

you're trying to find the difference in the volume now at 314 and when it is full

OpenStudy (thatonegirl_):

The formula for the volume of a cylinder is V=pi r^2 h. We will try to find the height for both. We can plug in some information into the formula that we already know.

OpenStudy (thatonegirl_):

Do you want to first find the height of the water when the cylinder if full or at the current volume?

OpenStudy (errinss):

\[(\pi)*5^2*h\]

OpenStudy (thatonegirl_):

Good! And now you can set this equal to one of the volumes. At full volume it is 471cm3 so you can add that to your equation for V.

OpenStudy (errinss):

\[471=(\pi)*5^2*h\]

OpenStudy (thatonegirl_):

Yup! The question said you could use 3.14 for pi. What will the simplified equation look like?

OpenStudy (errinss):

\[471=78.5*h\]

OpenStudy (thatonegirl_):

Mhm. And then get h (height) by itself by dividing both sides by 78.5

OpenStudy (errinss):

56.955414012738853503184713375796=H

OpenStudy (thatonegirl_):

How did you get that?

OpenStudy (errinss):

i divided both sides by 78.5 like you said to

OpenStudy (thatonegirl_):

471/78.5=6 and when you divide 78.5 by 78.5 that just equals 1 on that side

OpenStudy (thatonegirl_):

\[6=1\times h\]

OpenStudy (thatonegirl_):

\[6=h\]

OpenStudy (errinss):

oh lol huh i must've hit the wrong button

OpenStudy (thatonegirl_):

Lol no problem. So 6 is the height of the water when the cylinder is at full volume, 471cm3. Now try to find the height when the water is at 314cm3

OpenStudy (errinss):

\[341=(\pi)*25*H\]

OpenStudy (errinss):

\[341=78.5*h\]

OpenStudy (thatonegirl_):

Yup! And then solve for h, the height.

OpenStudy (errinss):

4.31

OpenStudy (thatonegirl_):

Wait a min. Is the volume 314 or 341?

OpenStudy (errinss):

314 lol whoops

OpenStudy (errinss):

4

OpenStudy (thatonegirl_):

Good! So now that we have the heights of the water when the cylinder is full and when it is somewhat filled, we can find the difference between the height. What is the difference between 6 and 4?

OpenStudy (errinss):

2

OpenStudy (thatonegirl_):

There you go!

OpenStudy (errinss):

THANK YOU!!!

OpenStudy (thatonegirl_):

You're welcome! :)

OpenStudy (thatonegirl_):

I hope it helped you!

OpenStudy (errinss):

yes it did thank you!!!